OpenStudy (anonymous):

(x+2)/(x^2+6x-7)=8/(x^2+3x-4)

OpenStudy (anonymous):

you want to solve for x? :)

OpenStudy (anonymous):

simplify? :)

OpenStudy (anonymous):

If it's simplification then factor first and you'll get:\[= [(x+2)/(x+7)(x-1)] - [8/(x-1)(x+4)]\] multiply by x +7 and x+4 to get the same denominator : \[= [(x+2)(x+4)-8(x+7)/(x-1)(x+7)(x+4)]\] \[= (x^2 -2x - 48)/(x-1)(x+7)(x+4)\] \[= (x-8)(x+6)/(x-1)(x+7)(x+4)\] That's it ^_^